OpenStudy (anonymous):

Find the constant sum for an ellipse with foci F1 (0, -8) and F2 (0, 8) and the point on the ellipse (0, 10).

OpenStudy (anonymous):

I know that the formula is something like this:\[d = \sqrt{(x _{2}-x_{1}) + (y_{2}-y_{1})}+ \sqrt{(x_{2}-x_{3}+(y_{2}-y_{3}}\] And then whenever you plug in the values you get \[\sqrt{(0-0)^2+(8+8)^2} + \sqrt{(0-0)^2 + (8-10)^2}\] And then from THAT you get \[\sqrt{0 + 16} + \sqrt{0+4}\] And then that simplifies to 4+ 2 which is 6... Did I do this right? lol